# John Mendonça1,2, Isabel Trancoso1,2    Alon Lavie3,4

### v2

@@ Abstract
As the capabilities of chatbots and their underlying LLMs continue to dramatically improve, evaluating their performance has increasingly become a major blocker to their further development. A major challenge is the available benchmarking datasets, which are largely static, outdated, and lacking in multilingual coverage, limiting their ability to capture subtle linguistic and cultural variations. This paper introduces  [Uncaptioned image] MEDAL, an automated multi-agent framework for generating, evaluating, and curating more representative and diverse open-domain dialogue evaluation benchmarks [^1] Code and data at github.com/johndmendonca/medal.. Our approach leverages several state-of-the-art LLMs to generate user-chatbot multilingual dialogues, conditioned on varied seed contexts. A strong LLM (GPT-4.1) is then used for a multidimensional analysis of the performance of the chatbots, uncovering noticeable cross-lingual performance differences. Guided by this large-scale evaluation, we curate a new meta-evaluation multilingual benchmark and human-annotate samples with nuanced quality judgments. This benchmark is then used to assess the ability of several reasoning and non-reasoning LLMs to act as evaluators of open-domain dialogues. We find that current LLMs struggle to detect nuanced issues, particularly those involving empathy and reasoning.
MEDAL: A Framework for Benchmarking LLMs as Multilingual Open-Domain Chatbots and Dialogue Evaluators
@@ 1 Introduction
The ubiquitous use of Large Language Models (LLMs) for the majority of NLP tasks has prompted the research community to largely shift its focus to their development (Minaee et al., 2025). As such, newer models are being introduced at an increasingly rapid pace and frequently supplant state-of-the-art results on several established benchmarks (Hendrycks et al., 2021; Li et al., 2024b). However, as these models rapidly advance, the capacity to evaluate their performance in meaningful ways has increasingly struggled to keep pace, becoming a significant impediment to the safe and effective adoption of new LLMs for tasks at large. As such, a critical reflection on whether the established benchmarks themselves continue to be effective and reliable measures of progress has become essential (Laskar et al., 2024).
@@
Figure 1: Overview of  Refer to caption MEDAL, with contributions highlighted in grey. ➊ Dialogue generation (§3): Generation of multi-turn, multilingual synthetic dialogues seeded from diverse contexts, between an LLM acting as a user (with utterances validated by a LLM-judge) and an LLM acting as a chatbot; ➋ Automated labelling (§4): Multidimensional evaluation of all dialogues by a strong LLM (GPT-4.1); ➌ Benchmark Curation (§5): Curation of a smaller meta-evaluation benchmark using human annotations.
@@
When examining open-domain dialogue evaluation within this rapidly evolving context, we observe significant deficiencies in the methodology used for meta-evaluation of LLM-Judges – LLMs acting as automated evaluators of chatbot performance. In particular, most existing benchmarks rely on obsolete chatbot responses conditioned on human-human interactions (Mendonça et al., 2024a), creating a mismatch with contemporary human-chatbot interaction paradigms. Furthermore, to the best of our knowledge, efforts to diversify benchmarks beyond English have been largely limited to translating existing English datasets into other desired languages, rather than natively curating data in these languages (Mendonca et al., 2023; Rodríguez-Cantelar et al., 2023), thus failing to capture potential linguistic and cultural nuances in dialogue quality. To address these gaps, this work presents an automated, multi-agent framework called  [Uncaptioned image]MEDAL [^2] Multilingual Evaluation of Dialogue-evaluators using Automated LLM-benchmarks.. MEDAL is designed to generate, evaluate, and curate more relevant and diverse dialogue evaluation benchmarks.
Our framework operates in three stages. First, we synthetically generate a large collection of dialogues in multiple languages, using several state-of-the-art LLMs acting as agents in the dialogue-generation process (§3). These LLMs act both as the chatbot and as the human user, and they are conditioned on diverse seed contexts designed to cover various axes of variability (topic, persona, affective state, and language). User-side generation is validated by an LLM-Judge to ensure coherence. By incorporating different LLMs as dialogue participants and varying seed contexts, we aim to minimize biases and diversify the quality of the generated dialogue dataset.
In the second stage, we use a state-of-the-art LLM (GPT-4.1) as a strong automated multi-dimensional evaluator of the large multilingual dialogue dataset generated in the first stage (§4). This provides a scalable method for assessing chatbot performance across various quality dimensions (e.g., coherence, commonsense, empathy, and language). The large-scale automated annotation uncovered significant performance differences across languages and LLM sizes, underscoring the need to expand dialogue evaluation beyond English.
Finally, informed by the large-scale automated evaluation and the dialogue source characteristics (i.e., the participating LLMs), we automatically curate a novel, meta-evaluation benchmark annotated by human experts (§5). The curation process aims to maximize the diversity of rated chatbots represented in the curated benchmark, and to ensure a balanced distribution of dialogue quality across multiple dimensions. One of the main strengths of MEDAL is that it is model agnostic and can be further extended to additional languages and/or LLMs with relative ease.
@@
Using the curated meta-evaluation benchmark, we conduct a comprehensive assessment of the ability of various state-of-the-art LLMs to perform as automated open-domain dialogue evaluators. Our analysis highlights several deficiencies in current models’ ability to accurately identify issues in dialogues, suggesting more nuanced affective and reasoning capabilities are required to achieve performance parity with human experts. Crucially, this methodology allows us, to the best of our knowledge, to perform the first systematic assessment comparing the open-domain dialogue evaluation capabilities of reasoning and non-reasoning models within a multilingual context. We summarise our contributions in Figure 1.
@@ 2.2 Meta-evaluation Benchmarks
While these and other recent benchmarks (detailed in Appendix A) offer valuable insights, they predominantly focus on English. A notable exception is the work by Rodríguez-Cantelar et al. (2023), which contains a subset of unique Chinese dialogues. However, only the English subset (60 dialogues, which were translated to Chinese and Spanish) contains chatbot responses obtained from LLM-based chatbots.
In light of these limitations, MEDAL positions itself as the first truly multilingual meta-evaluation benchmark that explicitly assesses the performance of a wide range of recent open-access LLMs across multiple languages.
@@ 3.1 Multi-agent Framework
Given the instruction following and creative writing abilities of LLMs, we prompt-instruct an LLM (acting as a human user and equipped with the topic) to begin and guide the conversation with a chatbot (another LLM, conditioned only on prior conversational context). Since our goal is to evaluate chatbots, we advocate that this approach constructs a realistic simulated representation of actual open-domain human-chatbot interactions  [^3] Although it is not the goal of this work, the generated dialogues could be used to finetune efficient chatbots.. Additional details, including our specific prompts, are described in Appendix C.
@@ 3.1.1 LLM as a User
To maximise diversity and minimise inherit biases, we condition the generation of dialogues using several contextual cues, including a scene description obtained from a commonsense knowledge graph, persona, an affective state and a target language. The inclusion of detailed information for contextualization purposes has been shown to outperform conversations sampled without context in terms of specificity and interestingness (Kim et al., 2023; Lin et al., 2024).
Scene description.
We use the templated sentence form of Atomic 10x (West et al., 2022) to describe scenes. Atomic 10x is a synthetically generated commonsense knowledge graph, consisting of symbolic triplets describing two events and the relation between them. Whenever the scene does not include an affective state (frequently found in relations of type xEffect or xReact), we randomly sample an affective state from a curated list of states, equally distributed across the quadrants of the circumplex model (Russell, 1980). The inclusion of an affective state in the scene description aims to reduce bias in casual dialogue generation, which in our initial experiments overwhelmingly presented tones of excitement and happiness.
Persona.
Since we want to have a simulation of human users interacting with chatbots, the inclusion of a persona provides a greater opportunity to diversify both the behaviour and style of the generated answers. We randomly sample personas from the publicly available subset of Persona Hub (Ge et al., 2024), a collection of diverse personas automatically curated from web data. Since all personas are gender neutral, we optionally sample a gender to reduce bias in gendered languages.
Language.
To address the lack of resources in other languages, previous work in the curation of multilingual dialogue datasets has translated existing English dialogues into other languages (Lin et al., 2021; Liu et al., 2023b). This approach to generating dialogues in non-English languages has serious limitations, the most obvious of which being the strong cultural biases introduced in the resulting translated dialogues (e.g., references to named entities, locations, and/or events). In our case, we keep the templated prompting sentence in English, but ask the LLM to act as a user from a specified country, and allow it to adapt the scene while generating directly in the specified native target language. This promotes better cultural diversity in contrast with translation [^4] A cherry-picked example from our dataset highlights this diversity for a dialogue discussing the music genre ”funk”: In Portuguese it typically refers to Funk carioca, whereas in English, it is associated with the 60s North American Funk..
First Turn Generation.
3.1.2 LLM as a Judge
@@
3.1.3 LLM as a Chatbot
We employ several open-access multilingual LLMs of different sizes as target chatbots. All models share the same system prompt, which conditions the models to generate natural, conversational language that is clear and easy to follow. In order to mimic other studies of open domain chatbot performance, the LLMs generating the chatbot responses have access to the dialogue history but are not explicitly provided with the external context (scene and persona) that was made available to the user agent.
@@ 3.2 Experiments
Other Hyperparameters.
We limit the feedback loop to 10 iterations in the first turn. This ensures dialogue generation for every context (scene-language-persona). However, for subsequent turns, we limit the loop to 5 iterations. Our analysis in a subset of dialogues found that 5 iterations was sufficient to ensure an average success rate of over 90%, and keeps costs below a factor of 4666Incorporating the feedback loop makes the process 4 times more expensive in terms of LLM queries, assuming response generation and evaluation steps have equal cost.. Dialogue length is capped at 10 turns due to cost constraints.
@@ 3.3 Results
In total, 38,400 dialogues were generated, stemming from 4,800 unique conversation starters (400 scenes, 2 user LLMs, and 6 languages). The automated filtering step then removed 6.44% of the dialogues, resulting in a final dataset of 35,927 instances. Statistics such as average turn length and diversity metrics are included in Appendix C.
Human Validation.
In order to evaluate the performance of the LLM as a user, we conduct a small scale human validation, where we ask annotators to rate the quality of the utterances generated by the LLM acting as a user in terms of their human-likeness, on a 1-5 Likert scale. Our annotator pool consists of a mix of graduate professionals in linguistics and computational linguistics, all with prior experience with annotations [^7] Finch et al. (2023a) reports low training pass rates when employing crowdsourcing platforms such as MTurk.. We assign 2 annotators per language, except for German and Spanish where we only employ a single annotator due to lack of annotators with professional level expertise in these languages. In total, 100 dialogues per language were evaluated. Even when selecting the lowest score, 98% of the dialogues were rated 4 or 5 (5:79.3%; 4:18.7%; 3:1.5%; 2:0.3%; 1:0%), with an average agreement (measured with Krippendorff’s $\alpha$) of .2592, denoting fair agreement.
4 Chatbot Evaluation
4.1 Automated Evaluation
We evaluate the chatbots used in our dialogue generation process (§3) with GPT-4.1. This automated evaluation consists of a multidimensional analysis of each chatbot’s responses throughout the dialogue using a subset of behaviours from the ABC-Eval framework (Finch et al., 2023a). More details regarding this automated analysis, including the full prompt used for instructing the LLM is available in Appendix D.
4.2 Results
Figure 2 presents the percentage of detected issues across the 8 dimensions (uninterpretable, unsafe, lacks Empathy, lacks Commonsense, repetitive, incoherent, irrelevant, and non-factual), together with the average overall score for all tested chatbots. Of the 8 assessed categories, "non-factual" appears to be the dimension in which even the stronger models struggle to some degree – the best performing model was found to expose non-factual information in 2.7% of its dialogues.
@@ 4.2 Results
The performance differences between models roughly align with model size, with the exception of Aya Expanse-8b, which is competitive in performance with its larger counterpart, Aya Expanse-32b and the 70B models. We observe noticeable differences in performance between these two groups for the majority of categories. However, for "unsafe", "repetitive" and "lacks Empathy", the performance differences are less pronounced across model sizes, which may indicate that this behaviour is not strongly associated with model size but with other factors such as post-training data.
@@
We observe that, in general, all models perform best in English. Additionally, the model’s "family" largely predicts its language performance. For example, Qwen models underperform in German, in particular in the "uninterpretable" category. A cursory check of some German dialogues for these models indicates that many of the responses contain Chinese or English words. We also note that Phi-3.5 and Llama-3.1-8B generally struggle more with Chinese. This shows that despite being advertised as multilingual, many of these models exhibit substantial performance differences, even across high-resource languages.
@@ 5 Meta-Evaluation Benchmark
In the final stage, we curate a smaller multilingual dialogue subset, specifically targeted to be used as a new benchmark for evaluating the performance of LLMs acting as automated dialogue evaluators. Despite the strong evaluation capabilities of LLMs, prior work has identified limitations when using LLMs as judges (Zheng et al., 2023; Wu and Aji, 2025). When used for dialogue evaluation, LLMs have been shown to struggle with coherence and commonsense reasoning (Finch et al., 2023b; Mendonça et al., 2024a). Figure 3 presents an example from our curated English dataset.
Language !Int !Safe !Emp !Com Rep !Coh !Rel !Fac Overall
EN 1.000 .6414 .4792 .5581 .2803 .5885 .4501 .7361 .4707
ZH .3258 .8236 .5415 .5811 .3981 .4236 .3220 .8347 .2622
FR .3716 .5811 .3877 .6581 .3877 .4212 .3886 .7967 .3508
PT .3337 .5522 .3926 .4243 .2908 .2419 .2562 .8415 .4182
Table 1: Inter-Annotator Agreement (IAA) for all dimensions across languages with 2 annotators.

### v3

@@ Abstract
Evaluating the quality of open-domain chatbots has become increasingly reliant on LLMs acting as automatic judges. However, existing meta-evaluation benchmarks are static, outdated, and lacking in multilingual coverage, limiting their ability to fully capture subtle weaknesses in evaluation. We introduce  [Uncaptioned image] MEDAL, an automated multi-agent framework for curating more representative and diverse open-domain dialogue evaluation benchmarks [^1] Code and data: github.com/johndmendonca/medal.. Our approach leverages several state-of-the-art LLMs to generate user-chatbot multilingual dialogues, conditioned on varied seed contexts. Then, a strong LLM (GPT-4.1) is used for a multidimensional analysis of the performance of the chatbots, uncovering noticeable cross-lingual performance differences. Guided by this large-scale evaluation, we curate a new meta-evaluation multilingual benchmark and human-annotate samples with nuanced quality judgments. This benchmark is then used to assess the ability of several reasoning and non-reasoning LLMs to act as evaluators of open-domain dialogues. Using MEDAL, we uncover that state-of-the-art judges fail to reliably detect nuanced issues such as lack of empathy, commonsense, or relevance.
MEDAL: A Framework for Benchmarking LLMs as Multilingual Open-Domain Dialogue Evaluators
@@ 1 Introduction
Figure 1: Overview of  Refer to caption MEDAL, a 3-step framework for curating multilingual meta-evaluation benchmarks.
@@
Dialogue generation (§3): Generation of multi-turn, multilingual synthetic dialogues seeded from diverse contexts, between an LLM acting as a user (with utterances validated by a LLM-judge) and the target chatbot; ➋ Automated labelling (§4): Multidimensional evaluation of all dialogues by a strong LLM (GPT-4.1); ➌ Benchmark Curation (§5): Curation of meta-evaluation benchmarks using human annotations used to assess LLM judges.
@@
As research increasingly focuses on Large Language Models (LLMs), new models are being developed at a pace that rapidly outstrips our ability to evaluate them meaningfully (Laskar et al., 2024). This growing gap between model capability and evaluation capacity has become a significant impediment to the safe and effective adoption of new LLMs for tasks at large. When examining open-domain dialogue evaluation within this rapidly evolving context, we observe significant deficiencies in the methodology used for benchmarking LLMs tasked with evaluating open-domain dialogues. In particular, most existing benchmarks rely on obsolete chatbot responses conditioned on human-human interactions (Mendonça et al., 2024a), creating a mismatch with contemporary human-chatbot interaction paradigms. Furthermore, to the best of our knowledge, efforts to diversify benchmarks beyond English have been largely limited to translating existing English datasets into other desired languages, rather than natively curating data in these languages (Mendonca et al., 2023; Zhang et al., 2023; Rodríguez-Cantelar et al., 2023), thus failing to capture potential linguistic and cultural nuances in dialogue quality.
To address these gaps, this work presents an automated, multi-agent framework called  [Uncaptioned image]MEDAL [^2] Multilingual Evaluation of Dialogue-evaluators using Automated LLM-benchmarks.. MEDAL is designed to generate, evaluate, and curate more relevant and diverse dialogue meta-evaluation benchmarks as follows:
  1. 1.
    Dialogue generation (§3): synthetically generate a large collection of dialogues in multiple languages, using several state-of-the-art LLMs acting as agents in the dialogue-generation process.
  2. 2.
    Automated labelling (§4): a strong LLM provides large-scale, multi-dimensional analysis of the generated dialogues, surfacing varying levels of failure across languages w.r.t empathy, commonsense, and even fluency.
  3. 3.
    Benchmark curation (§5): guided by these automated labels, we curate and human-annotate a balanced set of dialogues to form a meta-evaluation benchmark that directly tests LLM-judges.
@@
Our analysis highlights several deficiencies in current models’ ability to accurately identify issues in dialogues, suggesting more nuanced affective and reasoning capabilities are required to achieve performance parity with human experts. Crucially, this methodology allows us to perform the first systematic assessment comparing the open-domain dialogue evaluation capabilities of reasoning and non-reasoning models within a multilingual context.
In summary, our contributions are as follows:
  • We propose  [Uncaptioned image] MEDAL (Figure 1) for building multilingual meta-evaluation benchmarks, with modular generation, labelling, and curation stages [^3] While not the focus of this work, the large-scale raw dialogues and labels produced during MEDAL’s generation and automated annotation stages can be repurposed to pre-train lightweight evaluators or to fine-tune dialogue agents..
  • An automated analysis of chatbot responses shows that there are significant cross-lingual performance differences between different open-access LLMs.
  • A new meta-evaluation benchmark that explicitly measures the performance of LLMs as dialogue evaluators across six languages. Our analysis shows (1) a significant gap in performance between humans and LLMs, especially for more nuanced quality dimensions; and (2) reasoning models achieve better evaluative performance.
@@ 2.2 Meta-evaluation Benchmarks
While these and other recent benchmarks (detailed in Appendix A) offer valuable insights, they predominantly focus on English. A notable exception is the work by Rodríguez-Cantelar et al. (2023), which contains a subset of unique Chinese dialogues. Similarly, Zhang et al. (2023) extend coverage by translating many existing meta-evaluation benchmarks into multiple languages, though their focus remains on translating rather than designing new multilingual evaluation tasks. Additionally, only the English subset of Rodríguez-Cantelar et al. (2023) (60 dialogues, which were translated to Chinese and Spanish) contains chatbot responses obtained from LLM-based chatbots. In light of these limitations, MEDAL positions itself as the first truly multilingual meta-evaluation benchmark that explicitly assesses the performance of a wide range of recent open-access LLMs across languages.
@@ 3 Dialogue Generation
The first step in MEDAL is to synthetically generate dyadic dialogues between a human user (simulated by an LLM) and a chatbot. This scaffolding step provides three key advantages over curating static, human-chatbot dialogues: (1) scalability, allowing us to overcome expensive large-scale collection of native dialogues data across multiple languages; (2) contemporary relevance, enabling the on-demand generation of interactions that reflect modern conversational patterns and topics not found in older corpora; (3) error representation, as a large volume of dialogues allows us to better surface errors that become increasingly hard to find as chatbots improve.
To mitigate the risks of synthetic data, we adopt a multi-agent setup where we recruit several LLMs acting as the user (§3.1.1) and employ a response generation feedback loop using a LLM as a user judge (§3.1.2) which is externally validated by human annotators (§3.4).
@@ 3.1 Multi-agent Framework
Given the instruction following and creative writing abilities of LLMs for dialogue generation (Kim et al., 2023), we prompt-instruct an LLM (acting as a human user and equipped with the topic) to begin and guide the conversation with a chatbot (another LLM, conditioned only on prior conversational context).
@@ 3.1.1 LLM as a User
To maximise diversity and minimise inherit biases, we condition the generation of dialogues using several contextual cues, including a scene description obtained from a commonsense knowledge graph, persona, an affective state and a target language. The inclusion of detailed information for contextualization purposes has been shown to outperform conversations sampled without context in terms of specificity and interestingness (Kim et al., 2023; Lin et al., 2024). A detailed explanation of each contextual cue is presented in Appendix C.
3.1.2 LLM as a User Judge
@@
3.1.3 LLM as a Chatbot
We employ several open-access multilingual LLMs of different sizes as target chatbots. All models share the same system prompt, which conditions the models to generate natural, conversational language that is clear and easy to follow. In order to mimic other studies of open domain chatbot performance, the LLMs generating the chatbot responses have access to the dialogue history but are not explicitly provided with the external context (scene and persona) that was made available to the user.
3.2 Native Language Generation
To address the lack of resources in other languages, previous work in the curation of multilingual dialogue datasets has translated existing English dialogues into other languages (Lin et al., 2021; Liu et al., 2023). This approach to generating dialogues in non-English languages has several limitations. First, English-centric cultural artifacts like named entities, locations, and events remain unchanged after translation. More critically, translation does not offer a true, non-confounded evaluation of a model’s native multilingual capabilities.
In MEDAL, we keep the templated prompting sentence in English, but ask the LLM to act as a user from a specified country, and allow it to adapt the scene while generating directly in the native target language. This promotes better cultural diversity in contrast with translation [^4] A cherry-picked example from our dataset highlights this diversity for a dialogue discussing the music genre ”funk”: In Portuguese it typically refers to Funk carioca, whereas in English, it is associated with the 60s North American Funk.. Evidence that native generation yields more culturally appropriate dialogues is provided through a head-to-head comparison in Appendix C.4.
@@ 3.2 Native Language Generation
Label Definition
Uninterpretable (!Int) Outputs unclear or confusing text to the point that its meaning is difficult to determine.
Unsafe (!Safe) Promotes or suggests harmful actions, or is offensive or threatening.
Lacks Empathy (!Emp) Fails to express appropriate emotional understanding when the context calls for it.
Lacks Commonsense (!Com) Shows poor reasoning or contradicts basic, widely known everyday knowledge.
Repetitive (Rep) Unnecessarily repeats the same point, phrase, or idea in a way that adds no value.
Incoherent (!Coh) Contradicts itself or makes statements that do not logically follow from prior context.
Irrelevant (!Rel) Introduces unrelated content that does not align with the flow or subject of the conversation.
Non Factual (!Fac) Provides information that is objectively false or contradicts publicly verifiable facts.
Overall Overall assessment of the dialogue.
Table 1: Definition of issue labels used for the chatbot evaluation study and meta-evaluation benchmark.
@@ 3.3 Experiments
Statistics
In total, 38,400 dialogues were generated, stemming from 4,800 unique conversation starters (400 scenes, 2 user LLMs, and 6 languages). The automated filtering step then removed 6.44% of the dialogues, resulting in a final dataset of 35,927 instances. Additional statistics such as average turn length and diversity metrics are included in Appendix C.
3.4 Human Validation.
In order to evaluate the performance of the LLM as a user, we conduct a small scale human validation, where we ask annotators to rate the quality of the utterances generated by the LLM acting as a user in terms of their human-likeness, on a 1-5 Likert scale. Our annotator pool consists of a mix of graduate professionals in linguistics and computational linguistics, all with prior experience with annotations [^6] Finch et al. (2023a) reports low training pass rates when employing crowdsourcing platforms such as MTurk.. We assign 2 annotators per language, except for German and Spanish where we only employ a single annotator due to lack of annotators with professional level expertise in these languages. In total, 100 dialogues per language were evaluated. Even when selecting the lowest score among annotators, 98% of the dialogues were rated 4 or 5 (5:79.3%; 4:18.7%; 3:1.5%; 2:0.3%; 1:0%), with an average agreement (measured with Krippendorff’s $\alpha$) of .2592, denoting fair agreement.
4 Chatbot Evaluation
Given the scale of dialogues generated in Section 3, an exhaustive human annotation is prohibitorily expensive. As a solution, we conduct an automated labelling step to (i) surface a wide range of possibly infrequent conversational issues at scale, (ii) compare error distributions across model families, and (iii) construct a balanced and representative pool of dialogues for meta-evaluation benchmark curation.
4.1 Experimental Setup
We evaluate the performance of the chatbots with GPT-4.1. This automated evaluation consists of a multidimensional analysis of each chatbot’s responses throughout the dialogue using a subset of behaviours from the ABC-Eval framework (Finch et al., 2023a) identified in Table 1. More details regarding this automated analysis, including the full prompt used for instructing the LLM is available in Appendix D.
@@ 4.2 Results
Figure 2 presents the percentage of detected issues across the 8 dimensions (uninterpretable, unsafe, lacks empathy, lacks commonsense, repetitive, incoherent, irrelevant, and non-factual), together with the average overall score for all tested chatbots. Of the 8 assessed categories, non-factual appears to be the dimension in which even the stronger models struggle to some degree – the best performing model was found to output non-factual information in 2.7% of its dialogues.
@@
The performance differences between models roughly align with model size, with the exception of Aya Expanse-8b, which is competitive in performance with its larger counterpart, Aya Expanse-32b and the 70B models. We observe noticeable differences in performance between these two groups for the majority of categories. However, for unsafe, repetitive and non-factual, the performance differences are less pronounced across model sizes, which may indicate that this behaviour is not strongly associated with model size but with other factors such as post-training data.
@@
We observe that, in general, all models perform best in English. Additionally, the model’s "family" largely predicts its language performance. For example, Qwen models underperform in German, in particular in the uninterpretable category. A cursory check of some German dialogues for these models indicates that many of the responses contain Chinese or English words. We also note that Phi-3.5 and Llama-3.1-8B generally struggle more with Chinese. This shows that despite being advertised as multilingual, many of these models exhibit substantial performance differences, even across high-resource languages.
@@ 5 Meta-Evaluation Benchmark
While the resulting automated labelling of Section 4 provides scale, prior work has identified limitations when using LLMs as judges (Zheng et al., 2023; Wu and Aji, 2025). When used for dialogue evaluation, LLMs have been shown to struggle with coherence and commonsense reasoning (Finch et al., 2023b; Mendonça et al., 2024a). Given that the goal of MEDAL is to curate meaningful meta-evaluation benchmarks, we propose distilling this large pool of synthetically generated and automatically labelled dialogues into high-quality benchmarks annotated by professional annotators. Figure 3 presents an example from our benchmark together with an incorrect automated evaluation.